# Application of Laplace filters to the analysis of lattice time correlators

### v1

@@ V.2.1 Multistate hypothesis testing
A key observation in the discussion above is that eq.˜49 has no solution if the number of filters is smaller than the number of states, i.e., $r<N_{s}$. This statement can be reformulated statistically as a hypothesis rejection, which allows us to design tests for the number of statistically significant states in a given time range.
@@
As an example, we show in figs.˜5 and 6 representations of $p_{\Lambda}$ for two states $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$, and for various values of the initial time $t_{i}$. At early values of $t_{i}$, we observe that $H_{\Lambda}$ is systematically rejected except for a well-localized $p_{\Lambda}>\alpha$ region in the $(E_{0},E_{1})$ plane. In this region, the data can be described up to noise as the sum of two exponentials with well-constrained energies. As the initial time increases, we observe that the $p_{\Lambda}>\alpha$ region eventually does not constrain the location of $E_{1}$ anymore. This suggests, as expected, that for sufficiently large times the correlator can be statistically represented by a single exponential.
@@
where $T_{r,\mathrm{obs}}^{2}$ is the minimized $r$-state $T^{2}$ obtained from the data. In practice, such a minimum can be computed using standard minimization algorithms with eq.˜53. If one rejects $H_{r}$ with $p_{r}<\alpha$, it means that with $1-\alpha$ confidence level, the data cannot be described as a sum of $r$ exponentials on the chosen time range. However, if $H_{r}$ is not rejected, it does not necessarily imply that all $r$ exponentials contribute significantly to the correlator – it is possible that a subset of $\Lambda$ is sufficient. We now discuss an extended procedure to address this.
@@ V.2.2 Number of significant states
In step 1, it is important to note that the constraint $t_{i}>r_{\mathrm{max}}$ must hold to satisfy the requirements of eq.˜48. In step 2, one must be careful that as $r_{\mathrm{max}}$ increases, naive tests of the form $\mathrm{bar}{p}_{r}<\alpha$ will also increase the probability of erroneous rejections (family-wise error rate). This can be controlled at level $\alpha$ in several standard ways; in this work, we use the Holm-Bonferroni [23] method, which is a simple modification of the rejection threshold. It is important to notice that family-wise error rate control methods generally impose to decide in advance the number of hypothesis to be tested, which is part of the justification for step 1. Finally, if $H_{s}$ is rejected in step 3, we consider the procedure to be inconclusive. This can happen, for example, if $r_{\mathrm{max}}$ is too small and not enough states were added to obtain an acceptable description of the data.
In fig.˜7, we show the result of applying the procedure above to M0M-mes-ll, F1M-mes-lh, and F1M-mes-hh. The outcome is as expected: the number of significant states decreases as the initial time increases. Charmonium-like correlators are well known to generally suffer from considerable excited-state contamination, and we indeed observe from our representative data (F1M-mes-hh) that there are no initial times compatible with a single-state description. We find the procedure to be particularly promising for determining safe minimum times to attempt a single-state fit to the data. Indeed, if the number of significant states is equal to $1$, then the procedure above guarantees that the data are compatible with a single-state description at the $95\%$ confidence level. In the data presented in fig.˜7, the earliest safe initial time for single-state fits is $t_{i}=12$ for M0M-mes-ll, and $t_{i}=21$ for F1M-mes-lh. It is crucial to note that the proposed procedure is able to establish these results with no knowledge of the amplitudes $A_{n}$, which makes it considerably more robust than attempting a multi-exponential fit test on the data.
@@ V.2.3 Possible extensions
As part of the procedure proposed above, one also obtains the energy values for the significant states through the minimization of $T^{2}(\Lambda)$, opening the way for novel methods to determine energies from lattice data, without requiring multi-exponential model fits to the data. For example, we show in the right panel of fig.˜7 the energies of the significant states for the F1M-mes-hh obtained through the analysis described in the previous section. The values obtained are remarkably stable, and an encouraging sign that determining the spectrum through the response of applying Laplace filters is feasible. We will explore this further in future publications. In particular, one needs to understand how to rigorously define the statistical uncertainties of the energies obtained in fig.˜7. Additionally, this new method essentially determines the energies as the coefficients of a finite-difference equation canceling the correlator. This is also the mathematical foundation of the Prony method, and comparison with the outcome of solving a Prony GEVP is an interesting extension. However, due to the iterative nature of our method, we can already state that our method does not suffer from issues or ambiguities related to ordering eigenvalues, which is a known problem with GEVP-based methods.
@@ VI Application examples
where the amplitudes $A_{n}$ and energies $E_{n}$ are model parameters. As discussed in the previous section, the time dependence of an arbitrary $n$-point function is generally more complicated; however, it is common practice to analyze such functions according to a single time separation, and in an asymptotic time regime where the model above is a suitable description, up to negligible exponential corrections coming from higher-energy states and periodic contributions in time. Energies and amplitudes are then extracted by means of a $\chi^{2}$ minimization, i.e. by fitting some restricted time range of the data to eq.˜60 for a finite number of states $N$. As outlined in section˜II.2, this requires the inversion of the estimated correlation matrix between the data points that enter the fit. As discussed in section˜IV, applying Laplace filtering allows one to find a representation of the data that significantly reduces its CDR, making the numerical inversion of the correlation matrix more reliable. However, to be of practical use, we also require that the statistical uncertainties of extracted observables such as energies and amplitudes remain under control. Anticipating the results from section˜VI.1, we find that for the datasets under consideration here, the value of $\lambda_{0}$ that minimizes the CDR leads to larger statistical uncertainties on the extracted parameters $A_{n}$ and $E_{n}$. However, as can be inferred from figs.˜1 and 4, there is a wide range of values of $\lambda$ around $\lambda_{0}$ for which the CDR is reduced. In fact, based on the discussion about spectral aspects in the previous section, we will see that in practice there are additional benefits of Laplace filtering that can be exploited.
@@ VI.1 Mesonic 2-point functions
In fig.˜8 we show fit results obtained from a simple single-state (i.e. $N=1$ in eq.˜60) to the F1M-mes-lh dataset. We hold the maximum fit time $t_{f}=3N_{t}/8=36$ fixed and vary $t_{i}$ (shown on the $x$-axis). Making use of the symmetry present in this particular correlation function, all fits were performed to folded data, meaning that the correlation function has been symmetrized for time reversal (i.e. $C[t]\mapsto(C[t]+C[N_{t}-t])/2$, with the restriction $t\leq N_{t}/2$) prior to any subsequent operation. This is of practical convenience but not a necessity for the suggested method to work.
@@
The first type (green pluses) are fits to the unmodified data. We observe that the earliest fits with an acceptable $p$-value occur around $t_{i}\approx 20$, but the central values still shift until they reach a stable value around $t_{i}\approx 24$. This is independent of whether the fit is performed to the full or the thinned time range. We note that this is consistent with the determination of the number of significant states presented in fig.˜7 which found that the original correlation function can be described by a single state from $t_{i}\geq 21$.
@@
For all three types of fits, when additionally applying thinning (larger symbols), the CDR is further reduced and the $p$-values of the resulting fits are typically quantitatively better, but the fit results remain largely unchanged. For convenience, results from an extensive stability analysis entirely analogous to the one presented in Ref. [19] are superimposed and shown by the gray bands in the top two panels of fig.˜8. The choice of $\lambda=1.30$ was initially found by tuning $\lambda$ with the goal of lowering the $t_{i}$ for which acceptable fits first occur. However, we can interpret this value by noting that the stability analysis found $E_{1}=1.222(27)$, hence resulting in $\tilde{E_{1}}=1.29946$. Hence filtering with $\lambda=1.30$ effectively eliminates the first excited state from the correlation function. Furthermore, from fig.˜7 we then expect that after removing the first excited state, the filtered correlation function can be described by a single state from $t_{i}\approx 15$ in agreement with our findings.
@@
Figure 10: The same data as in fig.˜9. The unfiltered data is shown with open symbols, whilst the filtered data has been rescaled by the appropriate factor of $(\lambda_{D_{s}}^{2}-(\tilde{E}_{0}^{D_{s}})^{2})(\lambda_{\eta_{s}}^{2}-(\tilde{E}_{0}^{\eta_{s}})^{2})$, assuming $E_{0}^{D_{s}}=1.154$ and $E_{0}^{\eta_{s}}=0.4030$.
@@ VI.2 Application to semileptonic decays
In the following, we will explore the application of our method to more complicated processes. We start by considering the semileptonic decay of a $D_{s}$ meson to an $\eta_{s}$ meson, where a charm quark decays to a strange quark. For demonstration, we only consider the zero recoil (also known as $q^{2}_{\mathrm{max}}$) data point, where the initial and the final hadronic states are at rest. In addition to the $\eta_{s}$ and $D_{s}$ 2-point functions (see eq.˜60), the three point function $C^{cs\to ss}_{3}[t_{\mathrm{op}},t_{\mathrm{snk}}]$ contributes (where we have assumed that $t_{\mathrm{src}}=0<t_{\mathrm{op}}<t_{\mathrm{snk}}\leq N_{t}/2$). The spectral decomposition for this three-point function is
@@
In practice, the choice of $t_{\mathrm{snk}}$ is restricted by the temporal extent of the lattice, so one has to carefully investigate whether a region for $t_{\mathrm{op}}$ and $t_{\mathrm{snk}}$ can be found for which $R$ is approximately constant. The left-hand panel of fig.˜9 shows this ratio as a function of $t_{\mathrm{op}}$ for 5 choices of $t_{\mathrm{snk}}$. We find that even for the largest values of $t_{\mathrm{snk}}$, we cannot isolate a region where stability in $t_{\mathrm{op}}$ and $t_{\mathrm{snk}}$ is observed, indicating that for this dataset there is no region where a fit to a single state is sufficient. Following our observations from the previous sections, a natural question is whether applying suitably tuned Laplace filters to the data might allow for a transformation of the data in which the excited states are sufficiently reduced to identify a suitable time range where a single-state description of the data is appropriate.
Since the spectrum of the correlation functions remains unchanged, separate Laplace filters can be applied to any of the three correlation functions entering the construction of eq.˜64. For the sake of simplicity, and because we find this to be sufficient, we restrict ourselves to filtering the two-point functions that enter. The right-hand panel of fig.˜9 shows the Laplace-filtered version of eq.˜64 where, prior to constructing the ratio, the two-point functions $C_{2}^{cs}$ and $C_{2}^{ss}$ have been filtered using $\lambda^{2}_{D_{s}}=2.9$ and $\lambda^{2}_{\eta_{s}}=1.2$, respectively. We stress that the asymptotic values these ratios approach are different between the two panels. This is unsurprising since the former should asymptotically plateau to $M_{00}(A^{\eta_{s}}_{0}A^{D_{s}}_{0})^{-1/2}$, while for the latter we have the replacement $A^{P}_{0}\to{A_{\lambda_{P};0}^{P}}$ for $P=\eta_{s},D_{s}$ (see eq.˜61). However, from the same equation, the relationship between $A_{0}$ and $A_{\lambda;0}$ is analytically known once $E_{0}$ has been determined, so this does not pose any difficulty in practice. In fig.˜10, we jointly plot the original data (open symbols) together with the appropriately rescaled version of the ratio. For this we use ground state energy estimates from late times of the effective mass. Concretely, we use $E_{0}^{D_{s}}=1.154$ and $E_{0}^{\eta_{s}}=0.4030$. For a full comparison between the filtered and the unfiltered result, the value for the energies should be obtained from fits to the data and their uncertainties should be propagated. However, this is not expected to make a large difference for the particular dataset at hand since the statistical uncertainty on individual late-time effective mass data points is below 2 per-mil in both cases. From the plot we can draw several conclusions:
@@ VI.3 Application to neutral meson mixing
The final dataset we consider describes neutral meson mixing. There are several similarities between this and the case of semileptonic decays, but the required spectral analysis is simplified by the fact that the initial and final state have the same spectrum, so that eq.˜62 becomes
@@
Since the generation of this dataset does not require sequential sources and sources were placed on every second time plane (see Refs. [18, 19] for details), many choices of $t_{\mathrm{snk}}$ are computationally accessible, and our dataset comprises the values $t_{\mathrm{snk}}=16,18,\cdots,46$. The left-hand panel of fig.˜11 shows the effective bag parameter as a function of $t_{\mathrm{op}}-t_{\mathrm{snk}}/2$ for different values of $t_{\mathrm{snk}}$. For better readability, we do not display all values for $t_{\mathrm{snk}}$. Clearly, the data point $t_{\mathrm{op}}=t_{\mathrm{snk}}/2$ (plotted as 0 on the x-axis) is furthest away from both source and sink, and hence should have the smallest excited-state contamination. We observe that for this data point, very small values of $t_{\mathrm{snk}}$ are statistically incompatible with the larger values of $t_{\mathrm{snk}}$ and hence suffer from large excited-state contamination. However, these data points are statistically the most precise, since the uncertainties grow as $t_{\mathrm{snk}}$ is increased.
@@
so that Laplace filtering is likely to be usable to remove unwanted excited states. A comprehensive stability analysis analogous to that of Ref. [19] has been performed to determine the ground and first excited state parameters of the two- and three-point functions under consideration here, finding $E_{0}=0.94136(27)$, $E_{1}=1.190(21)$, and $B=0.8341(13)$. Drawing on that knowledge, we apply a Laplace filter to cancel the expected leading contamination. For the case of the mesonic two-point function that enters the normalization, we choose $\lambda^{2}_{\mathrm{mes}}=\tilde{E}_{1}^{2}$. For the three-point function, we note that the leading excited-state contamination in eq.˜67 is $E_{01}=(E_{0}+E_{1})/2$, and we therefore apply a filter with $\lambda^{2}_{\mathrm{mix}}=\tilde{E}_{01}^{2}$. We then construct the Laplace-filtered version of $B_{\mathrm{eff}}[t_{\mathrm{snk}}/2,t_{\mathrm{snk}}]$. Given that we have knowledge of the spectrum, we can also correct this value by the different normalization incurred from the Laplace filtering and hence present the original and the filtered data on the same scale. This is shown in the right-hand panel of fig.˜11. The black data points show the original data as a function of $t_{\mathrm{snk}}$. There is clear evidence for excited-state contamination for $t_{\mathrm{snk}}\lesssim 36$. For $t_{\mathrm{snk}}\gtrsim 36$ the individual values start to be noisier, which can be understood by referring to the left-hand plot. The red data points show the Laplace-filtered version of the data. As expected, we observe that already for $t_{\mathrm{snk}}\approx 26$ we obtain ground-state dominated data with similar precision to those of the earliest ground-state dominated data points of the unfiltered data. We further find that different choices of $t_{\mathrm{snk}}$ fluctuate more for the filtered data, but this is at least partially by virtue of the achieved decorrelation from using a Laplace filter and hence the data points carry more individual information.

### v2

@@ V.2.1 Multistate hypothesis testing
A key observation in the discussion above is that eq.˜50 has no solution if the number of filters is smaller than the number of states, i.e., $r<N_{s}$. This statement can be reformulated statistically as a hypothesis rejection, which allows us to design tests for the number of statistically significant states in a given time range.
@@
As an example, we show in figs.˜6 and 7 representations of $p_{\Lambda}$ for two states $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$, and for various values of the initial time $t_{i}$. At early values of $t_{i}$, we observe that $H_{\Lambda}$ is systematically rejected except for a well-localized $p_{\Lambda}>\alpha$ region in the $(E_{0},E_{1})$ plane. In this region, the data can be described up to noise as the sum of two exponentials with well-constrained energies. As the initial time increases, we observe that the $p_{\Lambda}>\alpha$ region eventually does not constrain the location of $E_{1}$ anymore. This suggests, as expected, that for sufficiently large times the correlator can be statistically represented by a single exponential.
@@
where $T_{r,\mathrm{obs}}^{2}$ is the minimized $r$-state $T^{2}$ obtained from the data. In practice, such a minimum can be computed using standard minimization algorithms with eq.˜54. If one rejects $H_{r}$ with $p_{r}<\alpha$, it means that with $1-\alpha$ confidence level, the data cannot be described as a sum of $r$ exponentials on the chosen time range. However, if $H_{r}$ is not rejected, it does not necessarily imply that all $r$ exponentials contribute significantly to the correlator – it is possible that a subset of $\Lambda$ is sufficient. We now discuss an extended procedure to address this.
@@ V.2.2 Number of significant states
In step 1, it is important to note that the constraint $t_{i}>r_{\mathrm{max}}$ must hold to satisfy the requirements of eq.˜49. In step 2, one must be careful that as $r_{\mathrm{max}}$ increases, naively conducting multiple tests of the form $\mathrm{bar}{p}_{r}<\alpha$ will increase the probability of erroneous rejections, effectively lowering the overall confidence level. This is a well-known issue in statistics generally referred to as family-wise error rate. A confidence level of $1-\alpha$ can be enforced using the Holm-Bonferroni [29] method, which lowers the rejection thresholds to
$\mathrm{bar}{p}_{r_{j}}<\frac{\alpha}{r_{\mathrm{max}}+1-j}\,$, (61)
where $j\in\{1,\dots,r_{\mathrm{max}}\}$ is such that the sequence $\mathrm{bar}{p}_{r_{j}}$ is ordered from the lowest to the highest $p$-value. It is important to notice that family-wise error rate control methods generally require deciding in advance the number of hypothesis to be tested, which is part of the justification for step 1. Finally, if $H_{s}$ is rejected in step 3, we consider the procedure to be inconclusive. This can happen, for example, if $r_{\mathrm{max}}$ is too small and not enough states were added to obtain an acceptable description of the data.
In fig.˜8, we show the result of applying the procedure above to M0M-mes-ll, F1M-mes-lh, and F1M-mes-hh. The outcome is as expected: the number of significant states decreases as the initial time increases. Charmonium-like correlators are well known to generally suffer from considerable excited-state contamination, and we indeed observe from our representative data (F1M-mes-hh) that there are no initial times compatible with a single-state description. We find the procedure to be particularly promising for determining safe minimum times to attempt a single-state fit to the data. Indeed, if the number of significant states is equal to $1$, then the procedure above guarantees that the data are compatible with a single-state description at the $95\%$ confidence level. In the data presented in fig.˜8, the earliest safe initial time for single-state fits is $t_{i}=12$ for M0M-mes-ll, and $t_{i}=21$ for F1M-mes-lh. It is crucial to note that the proposed procedure is able to establish these results with no knowledge of the amplitudes $A_{n}$, which makes it more attractive than attempting a non-linear multi-exponential fit test on the data.
@@ V.2.3 Possible extensions
As part of the procedure proposed above, one also obtains the energy values for the significant states through the minimization of $T^{2}(\Lambda)$, opening the way for novel methods to determine energies from lattice data, without requiring multi-exponential model fits to the data. For example, we show in the right panel of fig.˜8 the energies of the significant states for the F1M-mes-hh obtained through the analysis described in the previous section. The values obtained are remarkably stable, and an encouraging sign that determining the spectrum through the response of applying Laplace filters is feasible. We will explore this further in future publications. In particular, one needs to understand how to rigorously define the statistical uncertainties of the energies obtained in fig.˜8. Additionally, this new method essentially determines the energies as the coefficients of a finite-difference equation canceling the correlator. This is also the mathematical foundation of the Prony method, and comparison with the outcome of solving a Prony GEVP is an interesting extension. However, due to the iterative nature of our method, we can already state that our method does not suffer from issues or ambiguities related to ordering eigenvalues, which is a known problem with GEVP-based methods.
@@ VI Application examples
where the amplitudes $A_{n}$ and energies $E_{n}$ are model parameters. As discussed in the previous section, the time dependence of an arbitrary $n$-point function is generally more complicated; however, it is common practice to analyze such functions according to a single time separation, and in an asymptotic time regime where the model above is a suitable description, up to negligible exponential corrections coming from higher-energy states and periodic contributions in time. Energies and amplitudes are then extracted by means of a $\chi^{2}$ minimization, i.e. by fitting some restricted time range of the data to eq.˜62 for a finite number of states $N$. As outlined in section˜II.2, this requires the inversion of the estimated correlation matrix between the data points that enter the fit. As discussed in section˜IV, applying Laplace filtering allows one to find a representation of the data that significantly reduces its CDR, making the numerical inversion of the correlation matrix more reliable. However, to be of practical use, we also require that the statistical uncertainties of extracted observables such as energies and amplitudes remain under control. Anticipating the results from section˜VI.1, we find that for the datasets under consideration here, the value of $\lambda_{0}$ that minimizes the CDR leads to larger statistical uncertainties on the extracted parameters $A_{n}$ and $E_{n}$. However, as can be inferred from figs.˜1 and 4, there is a wide range of values of $\lambda$ around $\lambda_{0}$ for which the CDR is reduced. In fact, based on the discussion about spectral aspects in the previous section, we will see that in practice there are additional benefits of Laplace filtering that can be exploited.
@@ VI.1 Mesonic 2-point functions
In fig.˜9 we show fit results obtained from a simple single-state (i.e. $N=1$ in eq.˜62) fit to the F1M-mes-lh dataset. We hold the maximum fit time $t_{f}=3N_{t}/8=36$ fixed and vary $t_{i}$ (shown on the $x$-axis). Making use of the symmetry present in this particular correlation function, all fits were performed to folded data, meaning that the correlation function has been symmetrized for time reversal (i.e. $C[t]\mapsto(C[t]+C[N_{t}-t])/2$, with the restriction $t\leq N_{t}/2$) prior to any subsequent operation. This is of practical convenience but not a necessity for the suggested method to work.
@@
The first type (green pluses) are fits to the unmodified data. We observe that the earliest fits with an acceptable $p$-value occur around $t_{i}\approx 20$, but the central values still shift until they reach a stable value around $t_{i}\approx 24$. This is independent of whether the fit is performed to the full or the thinned time range. We note that this is consistent with the determination of the number of significant states presented in fig.˜8 which found that the original correlation function can be described by a single state from $t_{i}\geq 21$.
@@
For all three types of fits, when additionally applying thinning (larger symbols), the CDR is further reduced and the $p$-values of the resulting fits are typically quantitatively better, but the fit results remain largely unchanged. For convenience, results from an extensive stability analysis entirely analogous to the one presented in Ref. [25] are superimposed and shown by the gray bands in the top two panels of fig.˜9. The choice of $\lambda=1.30$ was initially found by tuning $\lambda$ with the goal of lowering the $t_{i}$ for which acceptable fits first occur. However, we can interpret this value by noting that the stability analysis found $E_{1}=1.222(27)$, hence resulting in $\tilde{E_{1}}=1.29946$. Hence filtering with $\lambda=1.30$ effectively eliminates the first excited state from the correlation function. Furthermore, from fig.˜8 we then expect that after removing the first excited state, the filtered correlation function can be described by a single state from $t_{i}\approx 15$ in agreement with our findings.
@@
Figure 11: The same data as in fig.˜10. The unfiltered data is shown with open symbols, whilst the filtered data has been rescaled by the appropriate factor of $(\lambda_{D_{s}}^{2}-(\tilde{E}_{0}^{D_{s}})^{2})(\lambda_{\eta_{s}}^{2}-(\tilde{E}_{0}^{\eta_{s}})^{2})$, assuming $E_{0}^{D_{s}}=1.154$ and $E_{0}^{\eta_{s}}=0.4030$.
@@ VI.2 Application to semileptonic decays
In the following, we will explore the application of our method to more complicated processes. We start by considering the semileptonic decay of a $D_{s}$ meson to an $\eta_{s}$ meson, where a charm quark decays to a strange quark. For demonstration, we only consider the zero recoil (also known as $q^{2}_{\mathrm{max}}$) data point, where the initial and the final hadronic states are at rest. In addition to the $\eta_{s}$ and $D_{s}$ 2-point functions (see eq.˜62), the three point function $C^{cs\to ss}_{3}[t_{\mathrm{op}},t_{\mathrm{snk}}]$ contributes (where we have assumed that $t_{\mathrm{src}}=0<t_{\mathrm{op}}<t_{\mathrm{snk}}\leq N_{t}/2$). The spectral decomposition for this three-point function is
@@
In practice, the choice of $t_{\mathrm{snk}}$ is restricted by the temporal extent of the lattice, so one has to carefully investigate whether a region for $t_{\mathrm{op}}$ and $t_{\mathrm{snk}}$ can be found for which $R$ is approximately constant. The left-hand panel of fig.˜10 shows this ratio as a function of $t_{\mathrm{op}}$ for 5 choices of $t_{\mathrm{snk}}$. We find that even for the largest values of $t_{\mathrm{snk}}$, we cannot isolate a region where stability in $t_{\mathrm{op}}$ and $t_{\mathrm{snk}}$ is observed, indicating that for this dataset there is no region where a fit to a single state is sufficient. Following our observations from the previous sections, a natural question is whether applying suitably tuned Laplace filters to the data might allow for a transformation of the data in which the excited states are sufficiently reduced to identify a suitable time range where a single-state description of the data is appropriate.
Since the spectrum of the correlation functions remains unchanged, separate Laplace filters can be applied to any of the three correlation functions entering the construction of eq.˜66. For the sake of simplicity, and because we find this to be sufficient, we restrict ourselves to filtering the two-point functions that enter. The right-hand panel of fig.˜10 shows the Laplace-filtered version of eq.˜66 where, prior to constructing the ratio, the two-point functions $C_{2}^{cs}$ and $C_{2}^{ss}$ have been filtered using $\lambda^{2}_{D_{s}}=2.9$ and $\lambda^{2}_{\eta_{s}}=1.2$, respectively. We stress that the asymptotic values these ratios approach are different between the two panels. This is unsurprising since the former should asymptotically plateau to $M_{00}(A^{\eta_{s}}_{0}A^{D_{s}}_{0})^{-1/2}$, while for the latter we have the replacement $A^{P}_{0}\to{A_{\lambda_{P};0}^{P}}$ for $P=\eta_{s},D_{s}$ (see eq.˜63). However, from the same equation, the relationship between $A_{0}$ and $A_{\lambda;0}$ is analytically known once $E_{0}$ has been determined, so this does not pose any difficulty in practice. In fig.˜11, we jointly plot the original data (open symbols) together with the appropriately rescaled version of the ratio. For this we use ground state energy estimates from late times of the effective mass. Concretely, we use $E_{0}^{D_{s}}=1.154$ and $E_{0}^{\eta_{s}}=0.4030$. For a full comparison between the filtered and the unfiltered result, the value for the energies should be obtained from fits to the data and their uncertainties should be propagated. However, this is not expected to make a large difference for the particular dataset at hand since the statistical uncertainty on individual late-time effective mass data points is below 2 per-mil in both cases. From the plot we can draw several conclusions:
@@ VI.3 Application to neutral meson mixing
The final dataset we consider describes neutral meson mixing. There are several similarities between this and the case of semileptonic decays, but the required spectral analysis is simplified by the fact that the initial and final state have the same spectrum, so that eq.˜64 becomes
@@
Since the generation of this dataset does not require sequential sources and sources were placed on every second time plane (see Refs. [24, 25] for details), many choices of $t_{\mathrm{snk}}$ are computationally accessible, and our dataset comprises the values $t_{\mathrm{snk}}=16,18,\cdots,46$. The left-hand panel of fig.˜12 shows the effective bag parameter as a function of $t_{\mathrm{op}}-t_{\mathrm{snk}}/2$ for different values of $t_{\mathrm{snk}}$. For better readability, we do not display all values for $t_{\mathrm{snk}}$. Clearly, the data point $t_{\mathrm{op}}=t_{\mathrm{snk}}/2$ (plotted as 0 on the x-axis) is furthest away from both source and sink, and hence should have the smallest excited-state contamination. We observe that for this data point, very small values of $t_{\mathrm{snk}}$ are statistically incompatible with the larger values of $t_{\mathrm{snk}}$ and hence suffer from large excited-state contamination. However, these data points are statistically the most precise, since the uncertainties grow as $t_{\mathrm{snk}}$ is increased.
@@
so that Laplace filtering is likely to be usable to remove unwanted excited states. A comprehensive stability analysis analogous to that of Ref. [25] has been performed to determine the ground and first excited state parameters of the two- and three-point functions under consideration here, finding $E_{0}=0.94136(27)$, $E_{1}=1.190(21)$, and $B=0.8341(13)$. Drawing on that knowledge, we apply a Laplace filter to cancel the expected leading contamination. For the case of the mesonic two-point function that enters the normalization, we choose $\lambda^{2}_{\mathrm{mes}}=\tilde{E}_{1}^{2}$. For the three-point function, we note that the leading excited-state contamination in eq.˜69 is $E_{01}=(E_{0}+E_{1})/2$, and we therefore apply a filter with $\lambda^{2}_{\mathrm{mix}}=\tilde{E}_{01}^{2}$. We then construct the Laplace-filtered version of $B_{\mathrm{eff}}[t_{\mathrm{snk}}/2,t_{\mathrm{snk}}]$. Given that we have knowledge of the spectrum, we can also correct this value by the different normalization incurred from the Laplace filtering and hence present the original and the filtered data on the same scale. This is shown in the right-hand panel of fig.˜12. The black data points show the original data as a function of $t_{\mathrm{snk}}$. There is clear evidence for excited-state contamination for $t_{\mathrm{snk}}\lesssim 36$. For $t_{\mathrm{snk}}\gtrsim 36$ the individual values start to be noisier, which can be understood by referring to the left-hand plot. The red data points show the Laplace-filtered version of the data. As expected, we observe that already for $t_{\mathrm{snk}}\approx 26$ we obtain ground-state dominated data with similar precision to those of the earliest ground-state dominated data points of the unfiltered data. We further find that different choices of $t_{\mathrm{snk}}$ fluctuate more for the filtered data, but this is at least partially by virtue of the achieved decorrelation from using a Laplace filter and hence the data points carry more individual information.